Suppose a single firm produces all of the output in a contestable market. The market inverse demand function is P = 200 -2Q, and the firm’s cost function is C(Q) = 8Q. Determine the firm’s equilibrium price and corresponding profits.

Price: $

Profits: $
Answer #1

Answer

A contestable market is the market where there is no barriers to entry and exit. If firms in a contestable market earns positive profit then new firm enters till each earns 0 Profit.

Thus In a contestable market each firm incurs 0 profit.

i.e. Profit = Total revenue - Total cost = 0

As, Total revenue = P*Q = Total Cost = C = 8Q

=> PQ = 8Q => P = 8.

Hence Price = $8

Profit = PQ - C = 8Q - 8Q = 0

Hence, Profit = $0.